OpenStudy (precal):

need help finding the second derivative of sinx+x^2y=10

OpenStudy (precal):

\[sinx+x^2y=10\]

myininaya (myininaya):

\[\cos(x)+x^2y'+2xy=0\] \[ x^2y'=-2xy-\cos(x)\] \[y'=\frac{-2xy-\cos(x)}{x^2}\]

myininaya (myininaya):

now apply quotient rule to find y''

OpenStudy (precal):

ok need a minute to draw it out. How do you get the fraction bar to work in equation editor?

myininaya (myininaya):

frac{ }{ }

OpenStudy (precal):

ok let me retype it again

OpenStudy (precal):

\[y''=\left[ -2x^2y-2x^3y'+x^2\sin(x)-4x^2y+2xcos(x) \right]/x^4\]

myininaya (myininaya):

\[y''=\frac{[-2(xy'+y)+\sin(x)]x^2-2x[-2xy-\cos(x)]}{x^4}\]

myininaya (myininaya):

i had to do it myself this one is a little more uglier

OpenStudy (precal):

yes it is. I wonder if I can put these into wolfram alpha?

myininaya (myininaya):

\[y''=\frac{-2x^3y'-2x^2y+x^2\sin(x)+4x^2y+2x \cos(x)}{x^4}\]

myininaya (myininaya):

yes i think you can i will look at it in a second and see okay?

OpenStudy (precal):

ok I see I made a sign mistake

myininaya (myininaya):

\[y''=\frac{-2x^3 (\frac{-2xy-\cos(x)}{x^2})-2x^2y+x^2 \sin(x)+4x^2y+2x \cos(x)}{x^4}\]

OpenStudy (precal):

are you using frac{}|{} in the equation editor ? I tried and it did not seem to work

myininaya (myininaya):

yes don't put that | thing in between the {} {}

myininaya (myininaya):

just put frac{ }{ }

myininaya (myininaya):

numerator inside first one denominator inside second one

OpenStudy (precal):

\[y'=\frac{2}{3}\]

OpenStudy (precal):

cool thanks, I just wrote any fraction to play around

OpenStudy (precal):

thanks for your help. I was playing around in wolframalpha and was not able to do the second derivatives.

myininaya (myininaya):

i'm still trying

myininaya (myininaya):

lol i don't know if i can figure it out you know how to use wofram to do it

myininaya (myininaya):

i can only get it to tell me the first derivative sorry

OpenStudy (precal):

myininaya can you tell me if the second derivative of \[2x^2+y^2=4\] is\[\frac{d^2y}{dx}=\frac{-8}{y^3}\]

OpenStudy (precal):

Thanks in advance for all of your help tonight :)

myininaya (myininaya):

\[4x+2yy'=0 => y'=\frac{-4x}{2y}=\frac{-2x}{y} => y''=\frac{-2y-(-2x)y'}{y^2}\] \[y''=\frac{-2y+2x \frac{-2x}{y}}{y^2}=\frac{-2y+-2x^2 \cdot \frac{1}{y}}{y^2} \cdot \frac{y}{y}\] \[y''=\frac{-2y^2-2x^2}{y^3}=\frac{-2(x^2+y^2)}{y^3}=\] .... i think i did something wrong

myininaya (myininaya):

oh i see it

OpenStudy (precal):

I got \[\frac{d^2y}{dx}=\frac{-4x^2-2y^2}{y^3}\]

myininaya (myininaya):

\[y''=\frac{-2y+2x \frac{-2x}{y}}{y^2}=\frac{-2y+-4x^2 \cdot \frac{1}{y}}{y^2} \cdot \frac{y}{y}\] \[y''=\frac{-2y^2-4x^2}{y^3}=\frac{-2(2x^2+y^2)}{y^3}=\]

myininaya (myininaya):

\[=\frac{-2(4)}{y^3}=\frac{-8}{y^3}\]

OpenStudy (precal):

Thanks you are awesome. I enjoy your company and help :)

myininaya (myininaya):

:)

OpenStudy (anonymous):

sinx+x^2 y=10 cosx + 2xy + x^2y' = 0 -> y' = (- cosx + 2xy ) / x^2 -> y" = x^2 [ sinx + 2( y + xy') ] - 2x (- cosx + 2xy ) / x^4 => y" = x^2sinx - 2x^2y + 2xcosx + 2x^3y' => y" = ( xsinx - 2xy +2cosx + 2x^2y') / x^3